HIGH RESOLUTION ELECTRONIC SPECTROSCOPY OF P-VINYLPHENOL IN THE GAS PHASE ^{a}

\underline{PHILIP J. MORGAN}, DIANE M. MITCHELL and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, PA 15260.

Recently, a controversy has developed over the proper assignment of the electronic spectrum of trans-p-coumaric acid (tPCA), the chromophore in photoactive yellow protein. Ryan et al.^{b} claim that two closely spaced peaks near 33,200 cm^{−1} are the S_{1} ← S_{0} origin bands of tPCA, whereas de Groot and Buma^{c} argue (based on REMPI results) that the spectrum should be contributed to the decomposition product p-vinylphenol (pVP). We have addressed this issue by recording the fully resolved spectra of these two bands. The derived values of the rotational constants show unambiguously that the carrier of these bands is pVP; the two conformers are due to the two possible orientations of the -OH group with respect to the vinyl group. With the aid of theoretical calculations, the origin at 33,207.3 cm^{−1} has been assigned to trans-pVP and the origin at 33,211.8 cm^{−1} to cis-pVP. These results confirm the thermal decarboxylation of tPCA.

aWork supported by NSF (CHE-0315584)

bW. L. Ryan, D. J. Gordon, and D. H. Levy, J. Am. Chem. Soc. 124, 6194 (2002).

cM. de Groot and W. J. Buma, J. Phys. Chem. A 109, 6135 (2005).